Species Lower Dorsey Stream Pig Pen Stream
Spotted dusky salamander 7 18
Imitator salamander 6 3
Seal salamander 5 15
Black-bellied salamander 7 11
Desmognathus spp. salamander 4 17
Blue Ridge two-lined salamander 1 31
Spring salamander 2 1
Northern slimy salamander 0 1
Black-chinned red salamander 0 0
Santeetlah salamander 1 0
Southern red-backed salamander 2 0

1. How was this set of species chosen for sampling? a. Taxonomic group b. Trophic group c. Guild d. Functional group

2. Determine the species richness (s) at each location and overall

3. Calculate species evenness using Shannon entropy (H) at each location and overall

4. Calculate alpha, gamma, and beta diversity (use species richness as the measure of diversity)
Answer #1

1. The answer would be a. Taxonomic group. All the samples belong to different species of the same genus Desmognathus.

2. Species richness is a measure of the sample, total how many species are present.

It is calculated as D = s/

where s= No of different species

N = Number of Individual organism

So, for Lower Dorsey Stream location, D = 9/35 = 1.52

and for Pig Pen Stream location D = 8/ 97 = 0.81

Overall = 10/ 132 = 0.87

3.  

Lower Dorsey Stream Shanon Index variable name Shanon Index calculation
Spotted dusky salamander 7 =n1 -0.32189
Imitator salamander 6 =n2 -0.30233
Seal salamander 5 =n3 -0.27799
Black-bellied salamander 7 =n4 -0.32189
Desmognathus spp. salamander 4 =n5 -0.24789
Blue Ridge two-lined salamander 1 =n6 -0.10158
Spring salamander 2 =n7 -0.16355
Northern slimy salamander 0 =n8 0
Black-chinned red salamander 0 =n9 0
Santeetlah salamander 1 =n10 -0.10158
Southern red-backed salamander 2 =n11 -0.16355
35 N 2.002255

This was done using the following formula for Shannon index - H’ = [(n1/N) ln(n1/N)]

Now Shannon's equitability (EH) or evenness can be calculated as

where S is the number of species.

SO EH = 2.002255/ln(9) = 2.098

Pig Pen Stream Shanon Index variable name Shanon Index calculation
Spotted dusky salamander 18 =n1 -0.31256
Imitator salamander 3 =n2 -0.10751
Seal salamander 15 =n3 -0.28866
Black-bellied salamander 11 =n4 -0.24686
Desmognathus spp. salamander 17 =n5 -0.30521
Blue Ridge two-lined salamander 31 =n6 -0.36456
Spring salamander 1 =n7 -0.04716
Northern slimy salamander 1 =n8 -0.04716
Black-chinned red salamander 0 =n9 0
Santeetlah salamander 0 =n10 0
Southern red-backed salamander 0 =n11 0
97 N 1.719676

So E H = 1.719676/ ln(8) = 1.904

Overall -

Total Shanon Index variable name Shanon Index calculation
Spotted dusky salamander 7 18 25 =n1 -0.31514
Imitator salamander 6 3 9 =n2 -0.18311
Seal salamander 5 15 20 =n3 -0.28592
Black-bellied salamander 7 11 18 =n4 -0.2717
Desmognathus spp. salamander 4 17 21 =n5 -0.29245
Blue Ridge two-lined salamander 1 31 32 =n6 -0.34353
Spring salamander 2 1 3 =n7 -0.086
Northern slimy salamander 0 1 1 =n8 0
Black-chinned red salamander 0 0 0 =n9 0
Santeetlah salamander 1 0 1 =n10 -0.03699
Southern red-backed salamander 2 0 2 =n11 -0.06348
35 97 132 N 1.878319

So E H = 1.878319/ln(10) = 1.878

4. alpha diversity for Lower Dorsey Stream is 9 and Pig Pen Stream is 8

Beta diversity = (No. of species sin area 1- the number of species common among both sites)+ (No. of species sin area 2- the number of species common among both sites)

=(9-7)+(8-7) = 2+1=3

Gamma diversity = Overall diversity of all areas = 10